# TiDES: The 4MOST Time Domain Extragalactic Survey

### v1

@@ Abstract
The Time Domain Extragalactic Survey (TiDES) conducted on the 4-metre Multi-Object Spectroscopic Telescope (4MOST) will perform spectroscopic follow-up of extragalactic transients discovered in the era of the NSF-DOE Vera C. Rubin Observatory. TiDES will conduct a 5-year survey, covering , and use around 250 000 fibre hours to address three main science goals: (i) spectroscopic observations of 30 000 live transients, (ii) comprehensive follow-up of 200 000 host galaxies to obtain redshift measurements, and (iii) repeat spectroscopic observations of Active Galactic Nuclei to enable reverberation mapping studies. The live spectra from TiDES will be used to reveal the diversity and astrophysics of both normal and exotic supernovae across the luminosity-timescale plane. The extensive host-galaxy redshift campaign will allow exploitation of the larger sample of supernovae and improve photometric classification, providing the largest-ever sample of spec-confirmed type Ia supernovae, capable of a sub-2 per cent measurement of the equation-of-state of dark energy. Finally, the TiDES reverberation mapping experiment of 700–1 000 AGN will complement the SN Ia sample and extend the Hubble diagram to .
keywords:
surveys – supernovae: general – galaxies: active – distance scale1 Introduction
@@
The NSF-DOE Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) heralds a new era of transient discovery: LSST will produce millions of transient alerts and photometric data for hundreds-of-thousands of supernovae (SNe) and other variable sources. While this discovery rate is exciting, the ability to obtain spectroscopic follow-up of these transients must similarly scale to enable much of the scientific exploitation. Follow-up spectra are critical for extracting the full astrophysical potential of the objects discovered, for example their classifications, rates, chemistry, redshifts, luminosities – and ultimately their physical nature. The capacity to assemble these spectroscopic samples presents a new challenge for the transient community, and to meet this, a next generation facility is needed.
@@
In this paper we present the Time Domain Extragalactic Survey (TiDES), which is designed to meet the spectroscopic requirements of the LSST-era for optical transients, their host galaxies, and active galactic nuclei (AGN). TiDES will use the 4-metre Multi-Object Spectroscopic Telescope (4MOST; de Jong et al., 2019), a fibre-fed spectroscopic survey facility on the European Southern Observatory (ESO) Visible and Infrared Survey Telescope for Astronomy (VISTA). This facility packs science fibres distributed over the 4.2 deg2 field-of-view, with each spine configurable by the AESOP fibre positioning system onto a target of interest. TiDES will exploit the fact that wherever 4MOST points over deg2 of the extragalactic sky, there will be known time-variable sources recently discovered by Rubin or other surveys. These will include recently discovered transients, and older, faded events for which host galaxy spectroscopy will be obtained for measuring redshifts. The strategy takes inspiration from the successful ‘OzDES’ survey (Yuan et al., 2015; Lidman et al., 2020), conducted using the 2dF fibre positioner and the AAOmega spectrograph on the Anglo-Australian Telescope, which secured live SN spectra and 4,500 redshifts (Vincenzi et al., 2019) for SN host galaxies over 27 deg2 of imaging from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016).
@@
TiDES will use around 250 000 fibre-hours to tackle three key science goals. The first is the use of type Ia supernovae (SNe Ia) as a cosmological probe. SNe Ia are an exceptionally mature and well-understood cosmological probe via their use as standardisable candles (Riess et al., 1998; Perlmutter et al., 1999; Betoule et al., 2014; DES Collaboration et al., 2024). They remain a uniquely powerful distance-indicator in the Universe, and directly constrain the properties of dark energy when combined with Planck Cosmic Microwave Background (CMB) (Planck Collaboration et al., 2019) and/or Baryon Acoustic Oscillation (BAO; e.g., DESI Collaboration et al., 2024) measurements. Current samples measure the average dark energy equation-of-state parameter to per cent (Scolnic et al., 2018; Brout et al., 2022; DES Collaboration et al., 2024) and show it to be broadly consistent with a cosmological constant (), with the intriguing possibility of a time-varying that motivates the need for larger samples of SNe Ia to higher-. When combined with SNe Ia detected by LSST, TiDES will enable the creation of a Hubble Diagram of cosmologically-useful SNe Ia and a per cent measurement of together with measurements of its time variation. Massive spectroscopic follow-up of SNe Ia will also allow their use in different contexts, for example their brightnesses are lensed by structure, tracing matter distributions along their line of sight (Refsdal, 1964; Holz, 1998; Metcalf, 1999; Jönsson et al., 2010; Smith et al., 2014; Macaulay et al., 2020; Shah et al., 2024b, a), and, with LSST, SNe Ia will be used to trace peculiar velocities on large scales and constrain the growth rate of structure (Howlett et al., 2017).
@@
The extragalactic time-domain universe is a far more diverse environment than imagined only a decade ago, and exploring the explosive timescale-luminosity phase-space forms the second goal of TiDES. Objects such as: superluminous supernovae (SLSNe; e.g., Quimby et al., 2011; Inserra et al., 2013; Angus et al., 2019; Kangas et al., 2022; Chen et al., 2023, for a review see Gal-Yam 2019), calcium-strong (-rich) transients (interchangeably referred to as CaST or CaRT in the literature e.g., Kasliwal et al., 2012; De et al., 2020), exotic thermonuclear explosions (see review by Taubenberger, 2017), fast, blue optical transients (FBOTs; e.g., Drout et al., 2014; Pursiainen et al., 2018; Prentice et al., 2018; Margutti et al., 2019; Ho et al., 2023), tidal disruption events (TDEs; e.g., Rees, 1988; Gezari et al., 2012; Holoien et al., 2014; van Velzen et al., 2021; Gezari, 2021), other ambiguous nuclear transients (Kankare et al., 2017; Trakhtenbrot et al., 2019; Wiseman et al., 2023), and kilonovae (Tanvir et al., 2013; Arcavi et al., 2017; Coulter et al., 2017; Lipunov et al., 2017; Smartt et al., 2017; Soares-Santos et al., 2017; Tanvir et al., 2017; Valenti et al., 2017; Levan et al., 2024, for a review see Margutti & Chornock 2021), have each demonstrated how little is known about explosive transient populations. TiDES will enlarge the population of spectroscopically-confirmed transients by an order of magnitude, and potentially uncover entirely new forms of explosions. This will include 9 000 core collapse SNe, 3 000 SLSNe for probing the universe, and 500 faint-and-fast SN Ia-like events for studying the full range of thermonuclear white dwarf explosions.
@@
The third goal of TiDES is cosmology and galaxy evolution with AGN. These are among the most energetic sources in the Universe, showing variability in all wavebands when mass is accreted onto supermassive black holes in the centres of galaxies. The variability of the optical continuum light from the accreting matter, together with its delayed response mirrored in the broad emission lines of the surrounding material can be used i) to dynamically measure black hole masses (e.g., Shen et al., 2016) and identify sub-parsec-scale supermassive black hole binaries among the AGN population, and ii) as a standardisable candle to high redshifts (Blandford & McKee, 1982; Peterson, 1993; Haas et al., 2011; Watson et al., 2011; Panda & Marziani, 2023; Czerny et al., 2023a). TiDES will establish an SNe-independent Hubble diagram of AGN and will deliver one of the, if not the, largest catalogue of dynamically measured supermassive black hole masses over redshifts . The expected accuracy of both the Hubble diagram as well as the black hole masses will depend critically on the achieved season lengths and total baseline of the survey (Malik et al., 2022) as well as characterisation of the objects in terms of luminosity and accretion rate (e.g. Du et al., 2016; Fonseca Alvarez et al., 2020; Prince et al., 2023; Zajaček et al., 2024a).
@@
This paper introduces the TiDES survey and demonstrates its potential using simulations of the LSST and 4MOST survey strategy. In Section 2 we describe the 4MOST spectrograph and the TiDES strategy. Section 3 details our TiDES simulations, and Section 4 presents sample statistics. Our expectation for the SN Ia cosmological analysis is detailed in Section 5. We then summarise in Section 6. Throughout, we assume a baseline cosmology with a matter density (e.g. Planck Collaboration et al., 2020) in a flat universe and a Hubble constant of km s-1 Mpc-1, and use the AB photometric system (Oke & Gunn, 1983).
@@
4MOST is a new high-multiplex, wide-field spectroscopic facility under development for VISTA, located on the ‘NTT peak’ at the Cerro Paranal in northern Chile. A full overview can be found in de Jong et al. (2019), with the scientific operations model described in Walcher et al. (2019), and the general survey strategy overview in Guiglion et al. (2019). 4MOST has a large field-of-view of 4.2 deg2 with a high multiplex capability: 1624 fibres feeding two low-resolution spectrographs (LRS), and 812 fibres feeding a high-resolution spectrograph (HRS). The LRS offers over 3700–9500Å, and the HRS offers in three 400–600Å wide passbands. Within a 1-hour observation, 4MOST has the sensitivity to obtain redshifts of mag galaxies or AGN using the LRS. TiDES will use the LRS fibres.
The 4MOST project as a whole consists of 25 consortium and community surveys each with their specific science goals. Each survey is characterised by a target selection algorithm, but all surveys share the focal plane and are observed in parallel, with individual targets for any given observation selected by a fibre-to-target assignment algorithm (Tempel et al., 2020b). It is expected that TiDES will average around 30–35 LRS fibres in each extragalactic 4MOST pointing – around two per cent of the LRS fibres available.
@@ 2.1.1 TiDES-Live
Work over the last decade has charted the luminosity–timescale plane for transients, and uncovered rare and new classes of transients. TiDES-Live will conduct a systematic campaign of follow-up that will be both unbiased and an order of magnitude larger in number than earlier efforts of this nature (e.g., Lidman et al., 2020), thereby providing improved rates for both new and well-understood phenomena with full photometric light curves. Early time spectra (3–4 days post-detection) for all classes of SNe will provide new insights into the progenitor environment and the outer layers of the ejecta, and, in 4MOST pointings with repeat visits, multiple spectra will trace SN evolution into the nebular phase. TiDES will also collect the largest spectroscopic sample of rare transients (e.g., calcium-strong fast transients and TDEs) to comprehensively study their important contributions to galactic chemical evolution.
For SN Ia cosmology, TiDES-Live will create an impressive Hubble diagram of more than 10 000 cosmologically-useful events over the redshift range for which there will be associated Rubin deep imaging. A further goal is to explore the dispersion in the Hubble diagram as a function of astrophysical properties (e.g., morphology, SN colour, SN spectral features), with the aim of improving the cosmological precision using SN Ia subsets. Moreover, as even the most advanced machine-learning SN photometric classification techniques are dependent on large, homogeneous and representative training samples (Lochner et al., 2016; Möller & de Boissière, 2020; Qu & Sako, 2022), the TiDES sample will become an unsurpassed training sample for SN photometric classifiers (Carrick et al., 2021) unlocking the potential of the entire LSST SN dataset.
@@ 2.1.2 TiDES-Hosts
In addition to providing galaxy redshifts essential for the LSST SN Ia Hubble Diagram, TiDES-Hosts will provide further diagnostics of the dispersion in SN Ia magnitudes, such as host galaxy metallicity and star-formation rate (Dumayne et al., 2023). Redshifts also enable the intrinsic properties of individual transients, like luminosity and evolution timescale, to be measured. TiDES-Hosts will provide the largest and most precise measure of the cosmological parameters over the key redshift range where dark energy becomes a dominant component.
@@ 2.1.3 TiDES-RM
The final survey concerns repeat spectroscopic observations of AGN for reverberation mapping (‘TiDES-RM’). The survey will focus on long-term variability monitoring of AGN in the deep drilling fields of LSST (Kovačević et al., 2022; Czerny et al., 2023b). Specifically, we will exploit AGN broad line lags to measure dynamical masses of supermassive black holes and construct a Hubble diagram out to redshift , i.e. into cosmic noon. In addition, the continuous high-quality spectroscopic monitoring will enable us to search for any unusual kinematics of the broad lines that could indicate binary supermassive black holes (for a review see De Rosa et al., 2019).
@@
As standardisable candles, AGN offer a complementary probe to SNe Ia with a reach to higher redshifts and with different astrophysical systematics. Recent work suggests that current probes at low and high redshifts may provide inconsistent results, beyond the well-known tension, pointing either towards unknown systematics in the current methods or new physics (e.g. Di Valentino et al., 2019; Cao et al., 2022; Khadka et al., 2023; Cao et al., 2024; Zajaček et al., 2024b). TiDES will move forward with a unique test for remaining systematic uncertainties by using two independent standardisable candles, thereby increasing the reliability of the final results. By targeting around 930 AGN over we can improve the precision of constraints based on reverberation mapped AGN by at least a factor of two (King et al., 2015).
@@
Of equal importance in TiDES-RM will be dynamical mass measures of large numbers of supermassive black holes (SMBH) to , i.e. into cosmic noon, the most active phase of galaxy and black hole growth. Supermassive black hole masses represent a key parameter in governing the assembly history of galaxies through associated feedback processes. Most current SMBH mass measurements outside the local Universe rely on indirect methods, most notably the single-epoch method that takes the width of a broad line and infers the mass via a set of calibrations against local scaling relations. This method is prone to biases and uncertainties as a single emission line width will be unable to capture the dynamical complexities of the underlying accretion and outflow processes and due to the intrinsic difference in AGN demographics at low and high redshift (e.g. Shankar et al., 2016; Woo et al., 2024). TiDES-RM will measure dynamical SMBH masses at redshifts up to via reverberation-mapping, which is the technique underlying the calibration of local scaling relations (e.g. Shen et al., 2024). In addition, the spatio-kinematic information from emission line lags can be used to trace outliers, with the potential of exposing binary supermassive black holes (e.g. Songsheng et al., 2019).

### v2

@@ Abstract
The Time Domain Extragalactic Survey (TiDES) conducted on the 4-metre Multi-Object Spectroscopic Telescope (4MOST) will perform spectroscopic follow-up of extragalactic transients discovered in the era of the NSF-DOE Vera C. Rubin Observatory. TiDES will conduct a 5-year survey, covering , and use around 250 000 fibre hours to address three main science goals: (i) spectroscopic observations of 30 000 live transients, (ii) comprehensive follow-up of 200 000 host galaxies to obtain redshift measurements, and (iii) repeat spectroscopic observations of Active Galactic Nuclei to enable reverberation mapping studies. The live spectra from TiDES will be used to reveal the diversity and astrophysics of both normal and exotic supernovae across the luminosity-timescale plane. The extensive host-galaxy redshift campaign will allow exploitation of the larger sample of supernovae and improve photometric classification, providing the largest-ever sample of type Ia supernovae, capable of a sub-2 per cent measurement of the equation-of-state of dark energy. Finally, the TiDES reverberation mapping experiment of 700–1 000 AGN will complement the SN Ia sample and extend the Hubble diagram to .
show]c.frohmaier@soton.ac.uk
1 Introduction
@@
The NSF-DOE Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ž. Ivezić et al., 2019) heralds a new era of transient discovery: LSST will produce millions of transient alerts and photometric data for hundreds-of-thousands of supernovae (SNe) and other variable sources. While this discovery rate is exciting, the ability to obtain spectroscopic follow-up of these transients must similarly scale to enable much of the scientific exploitation. Follow-up spectra are critical for extracting the full astrophysical potential of the objects discovered, for example their classifications, rates, chemistry, redshifts, luminosities – and ultimately their physical nature. The capacity to assemble these spectroscopic samples presents a new challenge for the transient community, and to meet this, a next generation facility is needed.
@@
In this paper we present the Time Domain Extragalactic Survey (TiDES), which is designed to meet the spectroscopic requirements of the LSST-era for optical transients, their host galaxies, and active galactic nuclei (AGN). TiDES will use the 4-metre Multi-Object Spectroscopic Telescope (4MOST; R. S. de Jong et al., 2019), a fibre-fed spectroscopic survey facility on the European Southern Observatory (ESO) Visible and Infrared Survey Telescope for Astronomy (VISTA). This facility packs science fibres distributed over the 4.2 deg2 field-of-view, with each spine configurable by the AESOP fibre positioning system onto a target of interest. Throughout this work we refer to the concept of a fibre-hour. A fibre-hour is one hour of observing time, including overheads, for one fibre. In total, 4MOST provides up to 2,436 fibre-hours per one hour of observation. TiDES will exploit the fact that wherever 4MOST points over deg2 of the extragalactic sky, there will be known time-variable sources recently discovered by Rubin or other surveys. These will include recently discovered transients, and older, faded events for which host galaxy spectroscopy will be obtained for measuring redshifts. The strategy takes inspiration from the successful ‘OzDES’ survey (F. Yuan et al., 2015; C. Lidman et al., 2020), conducted using the 2dF fibre positioner and the AAOmega spectrograph on the Anglo-Australian Telescope, which secured live SN spectra and 4,500 redshifts (M. Vincenzi et al., 2019) for SN host galaxies over 27 deg2 of imaging from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016).
@@
TiDES will use around 250 000 fibre-hours to tackle three key science goals. The first is the use of type Ia supernovae (SNe Ia) as a cosmological probe. SNe Ia are an exceptionally mature and well-understood cosmological probe via their use as standardisable candles (A. G. Riess et al., 1998; S. Perlmutter et al., 1999; M. Betoule et al., 2014; DES Collaboration et al., 2024). They remain a uniquely powerful distance-indicator in the Universe, and directly constrain the properties of dark energy when combined with Planck Cosmic Microwave Background (CMB) (Planck Collaboration et al., 2020) and/or Baryon Acoustic Oscillation (BAO; e.g., A. G. Adame et al., 2025) measurements. When combined with these probes (CMB, BAO), current SN samples measure the average dark energy equation-of-state parameter to per cent (D. M. Scolnic et al., 2018; D. Brout et al., 2022; DES Collaboration et al., 2024). This includes a systematic error budget that is comparable to the statistical uncertainty, and show to be broadly consistent with a cosmological constant (), with the intriguing possibility of a time-varying that motivates the need for larger samples of SNe Ia to higher-. Using SNe Ia detected by LSST, TiDES will enable the creation of a Hubble Diagram of cosmologically-useful SNe Ia and make a per cent measurement of , when including a CMB prior, and considering only statistical uncertainties. These constraints are a factor 10 improvement compared to DES-SN5YR. Massive spectroscopic follow-up of SNe Ia will also allow their use in different contexts, for example their brightnesses are lensed by structure, tracing matter distributions along their line of sight (S. Refsdal, 1964; D. E. Holz, 1998; R. B. Metcalf, 1999; J. Jönsson et al., 2010; M. Smith et al., 2014; E. Macaulay et al., 2020; P. Shah et al., 2024a, b), and, with LSST, SNe Ia will be used to trace peculiar velocities on large scales and constrain the growth rate of structure (C. Howlett et al., 2017).
@@
The extragalactic time-domain universe is a far more diverse environment than imagined only a decade ago, and exploring the explosive timescale-luminosity phase-space forms the second goal of TiDES. Objects such as: superluminous supernovae (SLSNe; e.g., R. M. Quimby et al., 2011; C. Inserra et al., 2013; C. R. Angus et al., 2019; T. Kangas et al., 2022; Z. H. Chen et al., 2023, for a review see A. Gal-Yam 2019), calcium-strong (-rich) transients (interchangeably referred to as CaST or CaRT in the literature e.g., M. M. Kasliwal et al., 2012; K. De et al., 2020), exotic thermonuclear explosions (see review by S. Taubenberger, 2017), fast, blue optical transients (FBOTs; e.g., M. R. Drout et al., 2014; M. Pursiainen et al., 2018; S. J. Prentice et al., 2018; R. Margutti et al., 2019; A. Y. Q. Ho et al., 2023), tidal disruption events (TDEs; e.g., M. J. Rees, 1988; S. Gezari et al., 2012; T. W. S. Holoien et al., 2014; S. van Velzen et al., 2021; S. Gezari, 2021), other ambiguous nuclear transients (E. Kankare et al., 2017; B. Trakhtenbrot et al., 2019; P. Wiseman et al., 2023), and kilonovae (N. R. Tanvir et al., 2013; I. Arcavi et al., 2017; D. A. Coulter et al., 2017; V. M. Lipunov et al., 2017; S. J. Smartt et al., 2017; M. Soares-Santos et al., 2017; N. R. Tanvir et al., 2017; S. Valenti et al., 2017; A. J. Levan et al., 2024, for a review see R. Margutti & R. Chornock 2021), have each demonstrated how little is known about explosive transient populations. TiDES will enlarge the population of spectroscopically-confirmed transients by an order of magnitude, and potentially uncover entirely new forms of explosions. This will include 9 000 core collapse SNe, 3 000 SLSNe for probing the universe, and 500 faint-and-fast SN Ia-like events for studying the full range of thermonuclear white dwarf explosions.
@@
The third goal of TiDES is cosmology and galaxy evolution with AGN. These are among the most energetic sources in the Universe, showing variability in all wavebands when mass is accreted onto supermassive black holes in the centres of galaxies. The variability of the optical continuum light from the accreting matter, together with its delayed response mirrored in the broad emission lines of the surrounding material can be used i) to dynamically measure black hole masses (e.g., Y. Shen et al., 2016) and identify sub-parsec-scale supermassive black hole binaries among the AGN population, and ii) as a standardisable candle to high redshifts (R. D. Blandford & C. F. McKee, 1982; B. M. Peterson, 1993; M. Haas et al., 2011; D. Watson et al., 2011; S. Panda & P. Marziani, 2023; B. Czerny et al., 2023a). TiDES will establish an SNe-independent Hubble diagram of AGN and will deliver one of the, if not the, largest catalogue of dynamically measured supermassive black hole masses over redshifts . The expected accuracy of both the Hubble diagram as well as the black hole masses will depend critically on the achieved season lengths and total baseline of the survey (U. Malik et al., 2022) as well as characterisation of the objects in terms of luminosity and accretion rate (e.g. P. Du et al., 2016; G. Fonseca Alvarez et al., 2020; R. Prince et al., 2023; M. Zajaček et al., 2024b).
@@
This paper introduces the TiDES survey and demonstrates its potential using simulations of the LSST and 4MOST survey strategy. In Section 2 we describe the 4MOST spectrograph and the TiDES strategy. Section 3 details our TiDES simulations, and Section 4 presents sample statistics. Our expectation for the SN Ia cosmological analysis is detailed in Section 5. We then summarise in Section 6. Throughout, we assume a baseline cosmology with a matter density (e.g. Planck Collaboration et al., 2020) in a flat universe and a Hubble constant of km s-1 Mpc-1, and use the AB photometric system (J. B. Oke & J. E. Gunn, 1983).
@@
4MOST is a new high-multiplex, wide-field spectroscopic facility under development for VISTA, located on the ‘NTT peak’ at the Cerro Paranal in northern Chile. A full overview can be found in R. S. de Jong et al. (2019), with the scientific operations model described in C. J. Walcher et al. (2019), and the general survey strategy overview in G. Guiglion et al. (2019). 4MOST has a large field-of-view of 4.2 deg2 with a high multiplex capability: 1624 fibres feeding two low-resolution spectrographs (LRS), and 812 fibres feeding a high-resolution spectrograph (HRS). The LRS offers over 3700–9500Å, and the HRS offers in three 400–600Å wide passbands. Within a 1-hour observation, 4MOST has the sensitivity to obtain redshifts of mag galaxies or AGN using the LRS. TiDES will use the LRS fibres.
The 4MOST project as a whole consists of 25 consortium and community surveys each with their specific science goals. Each survey is characterised by a target selection algorithm, but all surveys share the focal plane and are observed in parallel, with individual targets for any given observation selected by a fibre-to-target assignment algorithm (E. Tempel et al., 2020a). It is expected that TiDES will average around 30–35 LRS fibres in each extragalactic 4MOST pointing – around two per cent of the LRS fibres available.
@@ 2.1.1 TiDES-Live
Work over the last decade has charted the luminosity–timescale plane for transients, and uncovered rare and new classes of transients. TiDES-Live will conduct a systematic campaign of follow-up that will be both unbiased and an order of magnitude larger in number than earlier efforts of this nature (e.g., C. Lidman et al., 2020), thereby providing improved rates for both new and well-understood phenomena with full photometric light curves. Early time spectra (3–4 days post-detection) for all classes of SNe will provide new insights into the progenitor environment and the outer layers of the ejecta, and, in 4MOST pointings with repeat visits, multiple spectra will trace SN evolution into the nebular phase. TiDES will also collect the largest spectroscopic sample of rare transients (e.g., calcium-strong fast transients and TDEs) to comprehensively study their important contributions to galactic chemical evolution.
For SN Ia cosmology, TiDES-Live will create an impressive Hubble diagram of more than 10 000 cosmologically-useful events over the redshift range for which there will be associated Rubin deep imaging. A further goal is to explore the dispersion in the Hubble diagram as a function of astrophysical properties (e.g., morphology, SN colour, SN spectral features), with the aim of improving the cosmological precision using SN Ia subsets. Moreover, as even the most advanced machine-learning SN photometric classification techniques are dependent on large, homogeneous and representative training samples (M. Lochner et al., 2016; A. Möller & T. de Boissière, 2020; H. Qu & M. Sako, 2022), the TiDES sample will become an unsurpassed training sample for SN photometric classifiers (J. E. Carrick et al., 2021) unlocking the potential of the entire LSST SN dataset.
@@ 2.1.2 TiDES-Hosts
In addition to providing galaxy redshifts essential for the LSST SN Ia Hubble Diagram, TiDES-Hosts will provide further diagnostics of the dispersion in SN Ia magnitudes, such as host galaxy metallicity and star-formation rate (following techniques presented in e.g. J. Dumayne et al., 2023; L. Galbany et al., 2022). Redshifts also enable the intrinsic properties of individual transients, like luminosity and evolution timescale, to be measured. TiDES-Hosts will provide the largest and most precise measure of the cosmological parameters over the key redshift range where dark energy becomes a dominant component.
In this work we consider host galaxies from TiDES in isolation, but it is important to highlight the contribution of other surveys in 4MOST. There are several participating surveys such as the Wide-Area VISTA Extragalactic Survey (WAVES; S. P. Driver et al., 2019) and the 4MOST Hemisphere Survey of the Nearby Universe (4HS; E. N. Taylor et al., 2023) that will observe millions of galaxies over thousands-of-square-degrees. Naturally, some of these galaxies will host supernovae, this will benefit TiDES in two ways: i) TiDES will not need to spend fibre-hours re-observing these galaxies, ii) If 4MOST has completed its planned visits of any given field, redshifts will be available from galaxies hosting future supernovae. These additional observations will supplement the total number of available redshifts for any SN analysis.
@@ 2.1.3 TiDES-RM
The final survey concerns repeat spectroscopic observations of AGN for reverberation mapping (‘TiDES-RM’). The survey will focus on long-term variability monitoring of AGN in the deep drilling fields (DDFs) of LSST (A. B. Kovačević et al., 2022; B. Czerny et al., 2023b). Specifically, we will exploit AGN broad line lags to measure dynamical masses of supermassive black holes and construct a Hubble diagram out to redshift , i.e. into cosmic noon. In addition, the continuous high-quality spectroscopic monitoring will enable us to search for any unusual kinematics of the broad lines that could indicate binary supermassive black holes (for a review see A. De Rosa et al., 2019). 4MOST will routinely visit the DDFs every 10-14 days to observe the high density of requested targets and satisfy the deep observing requirements from other surveys. TiDES-RM will use each of the single epoch visits to sample the expected lags of our AGN at a similar rate to past reverberation mapping campaigns (e.g. A. J. Barth et al., 2011). The number of observations required for each object is determined by its lag length, with shorter lags being recovered first. Generally, for a successful recovery of the broad line lags, a single observing season should be longer than a lag length, and an entire observing campaign should be longer than 3 times the lag length.
@@
As standardisable candles, AGN offer a complementary probe to SNe Ia with a reach to higher redshifts and with different astrophysical systematics. Recent work suggests that current probes at low and high redshifts may provide inconsistent results, beyond the well-known tension, pointing either towards unknown systematics in the current methods or new physics (e.g. E. Di Valentino et al., 2019; S. Cao et al., 2022; N. Khadka et al., 2023; S. Cao et al., 2024; M. Zajaček et al., 2024a). TiDES will move forward with a unique test for remaining systematic uncertainties by using two independent standardisable candles, thereby increasing the reliability of the final results. By targeting around 930 AGN over we can improve the precision of constraints based on reverberation mapped AGN by at least a factor of two (A. L. King et al., 2015).
@@
Of equal importance in TiDES-RM will be dynamical mass measures of large numbers of supermassive black holes (SMBH) to , i.e. into cosmic noon, the most active phase of galaxy and black hole growth. Supermassive black hole masses represent a key parameter in governing the assembly history of galaxies through associated feedback processes. Most current SMBH mass measurements outside the local Universe rely on indirect methods, most notably the single-epoch method that takes the width of a broad line and infers the mass via a set of calibrations against local scaling relations. This method is prone to biases and uncertainties as a single emission line width will be unable to capture the dynamical complexities of the underlying accretion and outflow processes and due to the intrinsic difference in AGN demographics at low and high redshift (e.g. F. Shankar et al., 2016; J.-H. Woo et al., 2024). TiDES-RM will measure dynamical SMBH masses at redshifts up to via reverberation-mapping, which is the technique underlying the calibration of local scaling relations (e.g. Y. Shen et al., 2024). In addition, the spatio-kinematic information from emission line lags can be used to trace outliers, with the potential of exposing binary supermassive black holes (e.g. Y.-Y. Songsheng et al., 2019).